If Ptotal=PA+PB=Xa+P*a+Xb+P*b and Xa=1-Xb and Xb=1-Xa

How can Ptotal=(P*B-P*A)Xb-P*A ?
Answer #1

In the initial expression, Xa is replaced by 1 - Xb, leaving:

P total = (1 - Xb) P * A + Xb * P * B

Ptotal = P * A - P * A * Xb + Xb * PB

Common factor of Xb is taken out:

P total = P * A + Xb * (P * B - P * A)

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.